# Analysis and Simulation of a Fluid-Heat System in a Thin, Rough Layer in Contact With a Solid Bulk Domain

### v2

@@ 4 Homogenization
To pass to the limit $\varepsilon\to 0$ we utilize the concept of two-scale convergence for thin layers first introduced in [46, Definition 4.1] and further developed in following articles, see for example [14, 29] and the references therein. We start by introducing the definition, then consider the existence of limit functions, and then pass to the limit in the weak formulation. Lastly, we analyze the existence and uniqueness of solutions to the effective model.
@@ 4.1 Two scale convergence
Definiton 1 (Two-scale convergence on thin domains).
@@
    $\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}\int_{S}\int_{\Omega_{\varepsilon}}u_{\varepsilon}(t,x)\varphi\left(t,\tilde{x},\nicefrac{{x}}{{\varepsilon}}\right)\,\mathrm{d}x\,\mathrm{d}t=\int_{S}\int_{\Sigma}\int_{Y^{d}}u(t,\tilde{x},y)\varphi(t,\tilde{x},y)\,\mathrm{d}y\,\mathrm{d}\tilde{x}\,\mathrm{d}t,$ (4.1)
    for all $\varphi\in C(\overline{S\times\Sigma};C_{\#}(\overline{Y^{d}}))$. If in addition
@@
    $\lim_{\varepsilon\to 0}\int_{S}\int_{\Gamma_{\varepsilon}}u_{\varepsilon}(t,x)\varphi\left(t,\tilde{x},\nicefrac{{x}}{{\varepsilon}}\right)\,\mathrm{d}\sigma_{x}\,\mathrm{d}t=\int_{S}\int_{\Sigma}\int_{\Gamma}u(t,\tilde{x},y)\varphi(t,\tilde{x},y)\,\mathrm{d}\sigma_{y}\,\mathrm{d}\tilde{x}\,\mathrm{d}t,$ (4.2)
@@
Note, that due to the Assumption (A5) it holds for the boundary condition of the velocity
$\frac{1}{\sqrt{\varepsilon}}\|u_{\varepsilon,BC}\|_{L^{2}(\partial\Sigma\times(0,\varepsilon))}=\|u_{BC}\|_{L^{2}(\partial\Sigma\times(0,1))},$
as well as
$\begin{split}\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}\int_{\partial\Sigma\times(0,\varepsilon)}u_{\varepsilon,BC}\varphi\left(\tilde{x},\nicefrac{{x}}{{\varepsilon}}\right)\,\mathrm{d}x&=\lim_{\varepsilon\to 0}\int_{\partial\Sigma}\int_{0}^{1}u_{BC}(\tilde{x},x_{d})\varphi\left(\tilde{x},\nicefrac{{\tilde{x}}}{{\varepsilon}},x_{d}\right)\,\mathrm{d}x_{d}\,\mathrm{d}\tilde{x}\\ &=\int_{\partial\Sigma}\int_{Y^{d}}u_{BC}(\tilde{x},y_{d})\varphi(\tilde{x},y)\,\mathrm{d}y\,\mathrm{d}\tilde{x},\end{split}$ (4.3)
for all $\varphi\in C(\partial\Sigma;C_{\#}(\overline{Y^{d}}))$. The two above properties imply the strong two-scale convergence of the boundary function, $u_{\varepsilon,BC}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}u_{BC}$. Where (given that the underlying domain is rectangular with integer corners) the strong two-scale convergence on the lateral boundary is defined as Definition 1 $i)$, with $\Omega_{\varepsilon}$ and $\Sigma$ replaced by $\partial\Sigma\times(0,\varepsilon)$ and $\partial\Sigma$ respectively.
@@
Under the Assumption (A1)-(A5), it holds for the weak solution $(\theta_{\varepsilon},u_{\varepsilon},p_{\varepsilon})\in L^{2}(S;\Theta_{\varepsilon})\times V_{\varepsilon}\times Q_{\varepsilon}$ of Eq. 2.1 that there exists a subsequence still denoted by $(\theta_{\varepsilon},u_{\varepsilon},p_{\varepsilon})$ and:
@@
    a $\theta^{s}\in L^{2}(S;H^{1}(\Omega))\cap H^{1}(S;H^{1}(\Omega)^{\prime})$ with
    $\chi_{\Omega^{s}_{\varepsilon}}\theta_{\varepsilon}\to\theta^{s}\text{ in }L^{2}(S\times\Omega)\quad\text{ and }\quad\chi_{\Omega^{s}_{\varepsilon}}\nabla\theta_{\varepsilon}\rightharpoonup\nabla\theta^{s}\text{ in }L^{2}(S\times\Omega)^{d},$
@@
    a $\theta^{f}\in L^{2}(S;H^{1}(\Sigma))\cap H^{1}(S;H^{1}(\Sigma)^{\prime})$ and a $\theta_{1}^{f}\in L^{2}(S\times\Sigma;H^{1}_{\#}(Y^{d}))$ with
    $\chi_{\Omega^{f}_{\varepsilon}}\theta^{f}_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightarrow}}\chi_{Z}\theta^{f}\quad\text{ and }\quad\chi_{\Omega^{f}_{\varepsilon}}\nabla\theta^{f}_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}\left(\nabla_{\tilde{x}}\theta^{f}+\nabla_{y}\theta_{1}^{f}\right),$
@@
    a $u\in L^{2}(S\times\Sigma;H^{1}_{\#}(\overline{Z}))^{d}$ with
    $\chi_{\Omega^{f}_{\varepsilon}}u_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}u\quad\text{ and }\quad\chi_{\Omega^{f}_{\varepsilon}}\varepsilon\nabla u_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}\nabla_{y}u.$
    It holds $\operatorname{div}_{y}{(u(\tilde{x},y))}=0$ and $\operatorname{div}_{\tilde{x}}{\left(\int_{Z}u(\tilde{x},y)\,\mathrm{d}y\right)}=0$ for almost all $(\tilde{x},y)\in\Sigma\times Z$. Moreover, $u$ has, for almost all $(t,x)\in S\times\Sigma$, the boundary values $u(t,\tilde{x},\tilde{y},0)=u_{\text{motion}}$ for $\tilde{y}\in Y^{d-1}$ and $u(t,\tilde{x},y)=0$ for $y\in\Gamma$ almost everywhere.
@@
    a $p\in L^{2}_{0}(S\times\Sigma\times Z)$ that is periodic in $Y^{d-1}$ and
@@
$\lim_{\varepsilon\to 0}\int_{S}\int_{\Gamma_{\varepsilon}}\alpha(\theta^{s}_{\varepsilon}-\theta^{f}_{\varepsilon})\varphi\,\mathrm{d}\sigma_{x}\,\mathrm{d}t=\int_{S}\int_{\Sigma}\alpha|\Gamma|(\theta^{s}-\theta^{f})\varphi\,\mathrm{d}\tilde{x}\,\mathrm{d}t,$
for all admissible test functions $\varphi$.
@@
Given the estimates of Lemma 4 the convergence in $i)$ can be deduced with the same arguments as in [19, Theorem 4].
The aspects $ii)$-$iv)$ are obtained by the estimates of Lemma 4 and Lemma 7 and usual two-scale properties, see for example [14, 25, 29]. For the convergence of the time derivative $\partial_{t}\theta^{f}_{\varepsilon}$ as well as the strong two-scale convergence of $\theta^{f}_{\varepsilon}$, we point to the results of [27, Proposition 5], since we have the same $\varepsilon$-dependent solution estimates. There are already several results for fluid and pressure limits in thin layers, see for example [22, 26]. The general idea for the divergence properties is similar to the case of porous domains in non-thin domains [4, Proposition 1.14]. The last integral identity is also a common equality, see for example [46] for similar setups and the general proof idea. ∎
@@
Given the Assumptions (A2) it holds $\chi_{\Omega^{f}_{\varepsilon}}\mu(\theta^{f}_{\varepsilon})\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}\mu(\theta^{f})$ strongly in $L^{2}(S\times\Omega^{f}_{\varepsilon})$.
4.2 Homogenized equations
Given the previous results concerning the limit of the solution, we can now carry out the homogenization process in the weak formulation. The general procedure when determining the effective model consists of three steps:
  1. 1.
    Use smooth test functions, compatible with the two-scale concept, to pass to the limit $\varepsilon\to 0$.
  2. 2.
    Try to average and decouple the effective systems to simplify the equations and extract cell problems.
  3. 3.
    Apply a density argument to show that the limit also holds for test functions from more general spaces.
@@ 4.2 Homogenized equations
Most of the steps are standard for our case, therefore we do not explain everything in detail but only comment on the general behavior and explain the important terms more in-depth. The complete homogenized model is written down at the end of this section in Theorem 3.
Limit procedure for the heat system.
We start with temperature limit inside of $\Omega^{s}_{\varepsilon}$. Given the Assumptions (A6) and (A7) for the convergence of the initial condition and heat source, passing to the limit in the weak formulation can be carried out directly. To be exact, we do not require two-scale convergence inside the bulk domain. We point to [19, Section 5] for a similar setup with a rough interface.
Most of the equations for the temperature limit in $\Omega^{f}_{\varepsilon}$ are also quite standard and we point to [14, 46, 29] for similar setups. As usual for these kinds of diffusion problems, one obtains that the limit $\theta^{f}_{1}$ can be represented by
$\theta^{f}_{1}(t,\tilde{x},y)=\sum_{i=1}^{d-1}\omega_{i}(y)\partial_{x_{i}}\theta^{f}(t,\tilde{x})$
where $\omega_{i}\in H^{1}_{\#}(Z)$ is the weak solution of
$\begin{split}\operatorname{div}_{y}{(\nabla_{y}\omega_{i}+e_{i})}&=0\quad\text{ in }Z,\\ (\nabla_{y}\omega_{i}+e_{i})\cdot n&=0\quad\text{ on }\Gamma\cup\{y\in\partial Z:y_{d}=0\}.\end{split}$ (4.4)
Next, we have a closer look at the boundary values of the limit $\theta^{f}$. Take $\varphi\in C^{\infty}(\overline{S\times\Sigma})$ and $\psi\in C_{\#}^{\infty}(\overline{Z})^{d}$, such that $\operatorname{div}_{y}{\psi}=0$, $\psi(y)=0$, if $y\in\Gamma$ or $y_{d}=0$ and $\varphi_{|\Sigma^{\text{out}}}=0$. Testing with $\varphi(t,\tilde{x})\psi(\nicefrac{{x}}{{\varepsilon}})$ yields
$\frac{1}{\varepsilon}\int_{S}\int_{\Omega^{f}_{\varepsilon}}\theta^{f}_{\varepsilon}(t,x)\operatorname{div}_{x}{\left(\varphi(t,\tilde{x})\psi(\nicefrac{{x}}{{\varepsilon}})\right)}\,\mathrm{d}x\,\mathrm{d}t=-\frac{1}{\varepsilon}\int_{S}\int_{\Omega^{f}_{\varepsilon}}\nabla_{x}\theta^{f}_{\varepsilon}(t,x)\cdot\varphi(t,\tilde{x})\psi(\nicefrac{{x}}{{\varepsilon}})\,\mathrm{d}x\,\mathrm{d}t.$
Passing to the limit on both sides and afterward integrating by parts on the right side, we obtain
$\int_{S}\int_{\Sigma}\int_{Z}\theta^{f}(t,\tilde{x})\operatorname{div}_{\tilde{x}}{(\varphi(t,\tilde{x})\psi(y))}\,\mathrm{d}y\,\mathrm{d}\tilde{x}\,\mathrm{d}t=-\int_{S}\int_{\Sigma}\int_{Z}\left(\nabla_{\tilde{x}}\theta^{f}(t,\tilde{x})+\nabla_{y}\theta^{f}_{1}(t,\tilde{x},y)\right)\cdot\varphi(t,\tilde{x})\psi(y)\,\mathrm{d}y\,\mathrm{d}\tilde{x}\,\mathrm{d}t$
$=\int_{S}\int_{\Sigma}\int_{Z}\theta^{f}(t,\tilde{x})\operatorname{div}_{\tilde{x}}{(\varphi(t,\tilde{x})\psi(y))}\,\mathrm{d}y\,\mathrm{d}\tilde{x}\,\mathrm{d}t-\int_{S}\int_{\Sigma^{\text{in}}}\theta^{f}(t,\tilde{x})\varphi(t,\tilde{x})\int_{Z}\psi(y)\,\mathrm{d}y\cdot n_{\Sigma}\,\mathrm{d}\sigma_{\tilde{x}}\,\mathrm{d}t$
$+\int_{S}\int_{\Sigma}\int_{Z}\theta^{f}_{1}(t,\tilde{x},y)\varphi(t,\tilde{x})\operatorname{div}_{y}{(\psi(y))}\,\mathrm{d}y\,\mathrm{d}\tilde{x}\,\mathrm{d}t-\int_{S}\int_{\Sigma}\int_{\partial Z}\theta^{f}_{1}(t,\tilde{x},y)\varphi(t,\tilde{x})\psi(y)\cdot n_{Z}\,\mathrm{d}y\,\mathrm{d}\sigma_{\tilde{x}}\,\mathrm{d}t.$
Given the properties of $\psi$ and periodicity of $\theta^{f}_{1}$ the last two integrals are 0. So we obtain
$\int_{S}\int_{\Sigma^{\text{in}}}\theta^{f}(t,\tilde{x})\varphi(t,\tilde{x})\int_{Z}\psi(y)\,\mathrm{d}y\cdot n_{\Sigma}\,\mathrm{d}\sigma_{\tilde{x}}\,\mathrm{d}t=0.$
Since the above quantity has to hold for all $\psi$, we conclude that already
By a density argument we can conclude that this property has to hold for all $\varphi\in L^{2}(S;H^{1}_{\Sigma^{\text{out}}}(\Sigma))$, which yields that $\theta^{f}=0$ almost everywhere on $S\times\Sigma^{\text{in}}$. Note, that similarly the boundary values for $u$ on $\partial Z$ follow.
The other interesting part is the convection term. Here, Assumption (A5) as well as the strong two-scale convergence of $\theta^{f}_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightarrow}}\theta^{f}$ are used to pass to the limit. Take $\varphi_{\varepsilon}=\varphi_{0}(t,\tilde{x})+\varepsilon\varphi_{1}\left(t,\tilde{x},\nicefrac{{x}}{{\varepsilon}}\right)$ with $\varphi_{0}\in C^{\infty}(\overline{S\times\Sigma})$ and $\varphi_{1}\in C^{\infty}(\overline{S\times\Sigma},C_{\#}^{\infty}(\overline{Z}))$, such that $\varphi_{0}(t,\tilde{x})=0$ and $\varphi_{1}(t,\tilde{x},\nicefrac{{x}}{{\varepsilon}})=0$ if $\tilde{x}\in\Sigma^{\text{in}}$. In the case of the convection term, we can integrate by parts to first obtain
$\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}(u_{\varepsilon}\nabla\theta_{\varepsilon},\varphi_{\varepsilon})_{S\times\Omega^{f}_{\varepsilon}}$ $=\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}\Bigl{[}-(u_{\varepsilon}\theta_{\varepsilon},\nabla_{\tilde{x}}\varphi_{0})_{S\times\Omega^{f}_{\varepsilon}}-\varepsilon(u_{\varepsilon}\theta_{\varepsilon},\nabla_{\tilde{x}}\varphi_{1})_{S\times\Omega^{f}_{\varepsilon}}-(u_{\varepsilon}\theta_{\varepsilon},\nabla_{y}\varphi_{1})_{S\times\Omega^{f}_{\varepsilon}}$
$\hskip 42.67912pt+(u_{\varepsilon}\cdot n_{\varepsilon}\theta_{\varepsilon},\varphi_{0})_{S\times\partial\Omega^{f}_{\varepsilon}}+\varepsilon(u_{\varepsilon}\cdot n_{\varepsilon}\theta_{\varepsilon},\varphi_{1})_{S\times\partial\Omega^{f}_{\varepsilon}}\Bigr{]}$
Utilizing the strong two-scale convergence and (4.3) we can pass to the limit on the right side, producing
@@
$\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}(u_{\varepsilon}\nabla\theta_{\varepsilon},\varphi_{\varepsilon})_{S\times\Omega^{f}_{\varepsilon}}$ $=-(u\theta^{f},\nabla_{\tilde{x}}\varphi_{0})_{S\times\Sigma\times Z}-(u\theta^{f},\nabla_{y}\varphi_{1})_{S\times\Sigma\times Z}+(u_{BC}(\tilde{x},y_{d})\cdot n_{\Sigma}\theta^{f},\varphi_{0})_{S\times\partial\Sigma\times Y^{d}}.$
Integration by parts in the second term on the right-hand side, that $\operatorname{div}_{y}(u\theta^{f})=0$, the periodicity and boundary values of $u$ yield $(u\theta^{f},\nabla_{y}\varphi_{1})_{S\times\Sigma\times Z}=0$. Define $\bar{u}(t,\tilde{x})=\int_{Z}u(t,\tilde{x},y)\,\mathrm{d}y$. Then integrating the first term by parts yields

### v3

@@ 4 Homogenization
In this section, we investigate the limit behavior for $\varepsilon\to 0$ of the microscopic solution $(\theta_{\varepsilon},u_{\varepsilon},p_{\varepsilon})$ and identify the macroscopic limit problem which is solved by this limit (see Theorem 4). For this limit analysis, we utilize the concept of two-scale convergence for thin layers first introduced in [23, Definition 4.1] and further developed in following articles, see for example [38, 22] and the references therein. We start by introducing the definition, then consider the existence of limit functions, and then pass to the limit in the weak formulation. Lastly, we analyze the existence and uniqueness of solutions to the effective model.
@@ 4.1 Two scale convergence
Definition 1 (Two-scale convergence on thin domains).
@@
    $\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}\int_{S}\int_{\Omega_{\varepsilon}}u_{\varepsilon}(t,x)\varphi\left(t,\tilde{x},\nicefrac{{x}}{{\varepsilon}}\right)\,\mathrm{d}x\,\mathrm{d}t=\int_{S}\int_{\Sigma}\int_{Y^{d}}u(t,\tilde{x},y)\varphi(t,\tilde{x},y)\,\mathrm{d}y\,\mathrm{d}\tilde{x}\,\mathrm{d}t,$ (20)
    for all $\varphi\in C(\overline{S\times\Sigma};C_{\#}(Y^{d}))$. If in addition
@@
    $\lim_{\varepsilon\to 0}\int_{S}\int_{\Gamma_{\varepsilon}}u_{\varepsilon}(t,x)\varphi\left(t,\tilde{x},\nicefrac{{x}}{{\varepsilon}}\right)\,\mathrm{d}\sigma_{x}\,\mathrm{d}t=\int_{S}\int_{\Sigma}\int_{\Gamma}u(t,\tilde{x},y)\varphi(t,\tilde{x},y)\,\mathrm{d}\sigma_{y}\,\mathrm{d}\tilde{x}\,\mathrm{d}t,$ (21)
@@
We remark that Definition 1 is given for the thin layer $\Omega_{\varepsilon}=\Sigma\times(0,\varepsilon)$, since there the limit behavior is more straightforward to define. However, the definition can, with the extension operator from Lemma 1, be directly transferred to the rough thin layer $\Omega^{f}_{\varepsilon}$. A function $u_{\varepsilon}\in L^{2}(S\times\Omega^{f}_{\varepsilon})$ is said to two-scale converge to a function $u\in L^{2}(S\times\Sigma\times Y^{d})$ if
$\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}\int_{S}\int_{\Omega^{f}_{\varepsilon}}u_{\varepsilon}(t,x)\varphi\left(t,\tilde{x},\nicefrac{{x}}{{\varepsilon}}\right)\,\mathrm{d}x\,\mathrm{d}t=\int_{S}\int_{\Sigma}\int_{Z}u(t,\tilde{x},y)\varphi(t,\tilde{x},y)\,\mathrm{d}y\,\mathrm{d}\tilde{x}\,\mathrm{d}t,$
for all admissible test functions $\varphi$. In this case we write $u_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}u$.
@@
Under Assumptions (A1)-(A5), weak solutions $(\theta_{\varepsilon},u_{\varepsilon},p_{\varepsilon})\in L^{2}(S;\Theta_{\varepsilon})\times V_{\varepsilon}\times Q_{\varepsilon}$ of Eq. 1 have convergent subsequences. More specifically, there are limit functions (without relabeling for the subsequences)
@@
    $\theta^{s}\in L^{2}(S;H^{1}(\Omega))\cap H^{1}(S;H^{1}(\Omega)^{\prime})$ such that
    $\theta_{\varepsilon}^{s}\to\theta^{s}\text{ in }L^{2}(S\times\Omega),\quad\partial_{t}\theta_{\varepsilon}^{s}\rightharpoonup\partial_{t}\theta^{s}\text{ in }L^{2}(S;H^{1}(\Omega)^{\prime}),\quad\nabla\theta_{\varepsilon}^{s}\rightharpoonup\nabla\theta^{s}\text{ in }L^{2}(S\times\Omega)^{d},$
@@
    $\theta^{f}\in L^{2}(S;H^{1}(\Sigma))\cap H^{1}(S;H^{1}(\Sigma)^{\prime})$ and $\theta_{1}^{f}\in L^{2}(S\times\Sigma;H^{1}_{\#}(Y^{d}))$ such that
    $\theta^{f}_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightarrow}}\chi_{Z}\theta^{f}\quad\text{ and }\quad\nabla\theta^{f}_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}\left(\nabla_{\tilde{x}}\theta^{f}+\nabla_{y}\theta_{1}^{f}\right).$
    Further, it holds
    $\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}\int_{S}\langle\partial_{t}\theta_{\varepsilon}^{f},\psi+\varphi_{\varepsilon}\rangle_{H^{1}_{\#}(\Omega^{f}_{\varepsilon})}\,\mathrm{d}t=\int_{S}\int_{Z}\langle\partial_{t}\theta^{f},\psi+\varphi\rangle_{H^{1}_{\#}(\Sigma)}\,\mathrm{d}y\,\mathrm{d}t$
    for all $\eta\in L^{2}(S;H^{1}_{\#}(\Sigma))$ and $\varphi_{\varepsilon}(t,x)=\varphi(t,\tilde{x},\nicefrac{{x}}{{\varepsilon}})$ with $\varphi\in C_{0}^{\infty}(S;C_{\#}^{\infty}(\Sigma;C_{\#}(Z))).$
@@
    $u\in L^{2}(S\times\Sigma;H^{1}_{\#}(Y^{d}))^{d}$ with
    $u_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}u\quad\text{ and }\quad\varepsilon\nabla u_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}\nabla_{y}u.$
    It holds $\operatorname{div}_{y}{u(t,\tilde{x},y)}=0$ and $\operatorname{div}_{\tilde{x}}{\left(\int_{Z}u(t,\tilde{x},y)\,\mathrm{d}y\right)}=0$ for almost all $(t,\tilde{x},y)\in S\times\Sigma\times Z$. Moreover, $u=u_{\text{motion}}$ on $S\times\Sigma\times\{y_{d}=0\}$ and $u=0$ for $S\times\Sigma\times\Gamma$ almost everywhere.
@@
    $p\in L^{2}(S\times\Sigma;L^{2}_{0,\#}(Z))$ with
@@
$\lim_{\varepsilon\to 0}\int_{S}\int_{\Gamma_{\varepsilon}}\alpha(\theta^{s}_{\varepsilon}-\theta^{f}_{\varepsilon})\varphi(t,\tilde{x},\nicefrac{{x}}{{\varepsilon}})\,\mathrm{d}\sigma_{x}\,\mathrm{d}t=\int_{S}\int_{\Sigma}\alpha|\Gamma|(\theta^{s}-\theta^{f})\varphi(t,\tilde{x},y)\,\mathrm{d}\tilde{x}\,\mathrm{d}t,$ (22)
for all $\varphi\in C(\overline{S\times\Sigma};C_{\#}(\Gamma))$.
@@
$i).$ $\theta_{\varepsilon}^{s}$ is uniformly bounded in $L^{2}(S;H^{1}(\Omega^{s}_{\varepsilon}))\cap H^{1}(S;H^{1}(\Omega^{s}_{\varepsilon})^{\prime})$ as a result of lemmas 5 and 7. From here, these convergence results follow via [18, Theorem 4].
ii).ii). These convergence results are due to lemmas 5 and 7 together with the properties of two-scale convergence. For $\theta^{f}_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}\theta^{f}$ and $\nabla\theta^{f}_{\varepsilon}\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}\left(\nabla_{\tilde{x}}\theta^{f}+\nabla_{y}\theta_{1}^{f}\right)$, we refer to [22, Proposition 4.3]; the same arguments can also be found in [38, 48]. For the convergence of the time derivative $\partial_{t}\theta^{f}_{\varepsilon}$ as well as the strong two-scale convergence of $\theta^{f}_{\varepsilon}$, we point to [36, Proposition 5].
$iii)$-$iv.)$ There are already several very similar results for fluid and pressure limits in thin layers available in the literature, see for example [35, 49]. Here, we rely on the estimates given via Lemma 8. The specific convergence results in a stationary setting can be found in [35, Lemma 4.8] and can be easily transferred to our quasi-stationary setup due to the uniform $L^{2}$-bounds. The general proof idea for the divergence properties is the same as for porous media in non-thin domains [50, Proposition 1.14].
Finally, the last integral identity over interfaces (see Eq. 22) is also a common limit in similar setups, e.g., [38, 23, 36]. We observe that $\theta_{\varepsilon}^{f}$ and $\theta_{\varepsilon}^{s}$ are both uniformly bounded in $L^{2}(S\times\Gamma_{\varepsilon})$ via Lemma 5 together with Lemma 2. As a consequence, there are converging subsequences for which the limit Eq. 22 holds ([38, Theorem 4.4.(ii)(ii)] and [36, Lemma B.1, Corollary B.2]).
@@
Given the Assumption (A2) it holds $\chi_{\Omega^{f}_{\varepsilon}}\mu(\theta^{f}_{\varepsilon})\stackrel{{\scriptstyle 2}}{{\rightharpoonup}}\chi_{Z}\mu(\theta^{f})$ strongly in $L^{2}(S\times\Omega^{f}_{\varepsilon})$.
4.2 Homogenization limit
@@
We now carry out the homogenization process in the weak formulation. Some of the steps in this limit process are standard in this case; therefore we do not explain everything in detail but only comment on the general behavior and explain the important terms more in-depth.
Theorem 4 (Homogenized model).
Let the Assumptions (A1)-(A7) be fulfilled and let $(\theta^{s},\theta^{f},u,p)$ be the limit functions given by Proposition 3. Set $\bar{u}(t,\tilde{x})=\int_{Z}u(t,\tilde{x},y)\,\mathrm{d}y$. Then $(\theta^{s},\theta^{f},\bar{u},p)$ are weak solutions of the following problem:
$\partial_{t}\theta^{s}-\kappa^{s}\Delta\theta^{s}$ $=0\quad\quad$ $\text{ in }S\times\Omega,$ (23a)
$\theta^{s}(0,\cdot)$ $=\theta^{s}_{0}$ $\text{ in }\Omega,$ (23b)
@@
$-\kappa^{s}\nabla\theta^{s}\cdot\nu$ $=0$ $\text{ on }S\times\partial\Omega\setminus\Sigma,$ (23c)
$-\kappa^{s}\nabla\theta^{s}\cdot\nu$ $=\alpha|\Gamma|\left(\theta^{f}-\theta^{s}\right)+\bar{f}^{s}\quad$ $\text{ on }S\times\Sigma,$ (23d)
where $\bar{f}^{f,s}(t,\tilde{x})\coloneqq\int_{\Gamma}f^{f,s}(t,\tilde{x},y)\,\mathrm{d}\sigma_{y}$. The fluid temperature $\theta^{f}$ satisfies